My Birthday is May 16th 1985. My mom's Birthday is March 10th 1957. What day will she be twice as old as me?
Answer by richwmiller(17219)

18 years, 2 months, 6 days
6642 days
March 10th 2014 to May 16th 2014
67 days
March 10th 1957 to 7/5/ 2014
20936 days
May 17th 1985 to 7/5/ 2014
10642 days
(10642+x)*2=20936+x
21284-20936+2x=x

348+x=0
she was twice as old as you 348 days ago on 7/22/2013
20588 days and 10294 days old